a titration of potassium hydrogen phthalate required 38.14 mL of NaOH solution to reach an end point detected by phenolphthalen indicator. find the concentration of naoh


1
Expert's answer
2021-05-26T07:07:45-0400

"Molarity = moles\/litre"

Moles= 38.14/40= 0.9535

Molarity= 0.9535/1000

Molarity= 0.9× 10–³M